A firm is producing 4 units of output at an average total cost of $50. When the firm produces 5 units of output, average total cost falls to $45. What is the marginal cost of the fifth unit of output?

A. $10

B. $25

C. $45

D. $70
Answer #1

Answer: B. $25

Total cost for 4 units of output = 50*4 = $200

Total cost for 5 units of output = 45*5 = $225

Marginal cost of the fifth unit of output = $225-$200 = $25